OpenStudy (anonymous):

The recommended daily allowance (RDA) of the trace metal magnesium is 410\(rm mg/day\) Express this quantuty in \(\rm\mu g/day\)

OpenStudy (anonymous):

\(1 kg = 10^3 mg = 10^6 \mu g = 10^9 ng\) As you can see, \(1mg=10^3 \mu g\) so \(410mg=410 \times 10^3 \mu g\)